# Few-Shot Anomaly Detection via Category-Agnostic Registration Learning

### v1

@@ Abstract
Most existing anomaly detection methods require a dedicated model for each category. Such a paradigm, despite its promising results, is computationally expensive and inefficient, thereby failing to meet the requirements for real-world applications. Inspired by how humans detect anomalies, by comparing a query image to known normal ones, this paper proposes a novel few-shot anomaly detection (FSAD) framework. Using a training set of normal images from various categories, registration, aiming to align normal images of the same categories, is leveraged as the proxy task for self-supervised category-agnostic representation learning. At test time, an image and its corresponding support set, consisting of a few normal images from the same category, are supplied, and anomalies are identified by comparing the registered features of the test image to its corresponding support image features. Such a setup enables the model to generalize to novel test categories. It is, to our best knowledge, the first FSAD method that requires no model fine-tuning for novel categories: enabling a single model to be applied to all categories. Extensive experiments demonstrate the effectiveness of the proposed method. Particularly, it improves the current state-of-the-art for FSAD by 11.3% and 8.3% on the MVTec and MPDD benchmarks, respectively. The source code is available at https://github.com/Haoyan-Guan/CAReg.
@@
Anomaly detection, few-shot learning, self-supervised learning, registration.
@@ I Introduction
Figure 1: (a) The one-model-per-category paradigm for the vanilla anomaly detection and few-shot anomaly detection. (b) The one-model-all-category paradigm for the proposed category-agnostic few-shot anomaly detection.
@@
Inspired by how human beings detect anomalies, this paper explores a one-model-all-category paradigm for FSAD, i.e., training a category-agnostic AD model. To find the anomalies in an image, humans simply compare it to a normal exemplar and identify any difference. This approach requires no knowledge of the image category because comparison is naturally category agnostic. We thus model anomaly detection as a comparison task to make it category agnostic. To deal with the fact that objects may appear in different positions, orientations, or poses, the images are first transformed into one coordinate system through a form of registration [17, 18, 19] to facilitate the comparison.
An overview of the proposed Category-Agnostic Registration (CAReg) framework for FSAD is shown in Figure 1 (b). A Siamese network [20] containing three spatial transformer network blocks [21] is employed for registration. For better robustness, we perform registration at the feature level, instead of pixel-by-pixel image registration [19]. The feature-level registration loss, a relaxed version of the image-level pixel-wise registration loss, is used to minimize the cosine distance of the features from images of the same category. The trained model has the ability to perform registration across different categories and is able to be used at test-time directly on novel categories. Anomalies are detected by comparing the registered features of the novel test category image and its corresponding support set (a few normal images from the target category). The comparison process first estimates the normal distribution based on the support set, which can be implemented with statistical-based normal distribution estimators [22], and then fits the test samples to the normal distribution. Samples that fall out of the normal distribution are considered anomalies.
@@
  • Feature registration is leveraged as the proxy task to train anomaly detection models, which demonstrate increased generalizability across different categories and datasets.
  • CAReg outperforms the state-of-the-art FSAD methods on both anomaly detection and anomaly localization tasks with standard benchmark datasets.
@@
The rest of the paper is organized as follows. Sec. II reviews existing works related to AD and few-shot learning. Sec. III formulates the problem of FSAD. Sec. IV describes the category-agnostic registration training used for CAReg. Sec. V describes an AD system combining CAReg with multiple normal distribution estimators. Sec. VI presents the experimental results on several real-world datasets, showing the effectiveness of our method in both few-shot anomaly detection and anomaly localization. Finally, conclusions are drawn in Sec. VII.
@@ II-A Anomaly Detection
Anomaly detection (AD) [28] has a wide range of applications, encompassing still image AD [9, 10, 7, 8] and anomalous human behavior detection [29, 30, 31, 32, 33, 34]. The focus of this paper is on industrial defect detection [23, 24], covering both image-level anomaly detection and pixel-level anomaly localization. The main challenge of unsupervised AD, compared to supervised approaches, lies in the absence of training annotations, be it at the image-level and pixel-level. Models are expected to be trained solely with normal instances, excluding anomalies. Unsupervised AD research can be broadly categorized into two tracks: one-class classification based AD approaches and self-supervised learning-based AD approaches.
@@
Recent advances in one-class classification for AD increasingly rely on feature embeddings from pre-trained networks, leveraging these as robust indicators of normalcy. Various statistical estimators of normal distributions [22, 25, 26] help pinpoint significant deviations in feature behavior, flagging them as anomalies. Measures like the Mahalanobis distance [22] are frequently used to calculate anomaly scores. Techniques such as those in [22, 26, 42] have demonstrated notable success by utilizing models pretrained on the ImageNet dataset [43]. However, solely relying on ImageNet pre-training has its limitations. This paper introduces feature registration as a novel proxy task that enhances feature aggregation and generalization capabilities, thus refining anomaly detection performance.
@@ II-B Few-Shot Learning
Few-shot learning (FSL) addresses the task of adapting a model to novel categories with limited available images. FSL methods can be broadly categorized into metric learning [64, 65, 66], generation [67, 68, 69], and optimization [70, 71, 72] approaches. Metric learning methods aim to structure the feature space to find the closest distance between test images and known classes. Generation methods involve designing generative models to produce novel class images, thereby bridging the few-shot gap. Optimization methods identify commonalities among different classes and use optimization models for novel classes based on these commonalities. In this paper, CAReg is based on metric learning to address the challenge of FSAD. The challenge is to detect image anomalies for novel categories that are unseen in the training dataset, using only a few normal samples. The key point is to enable the model to be category-agnostic, and this ability is provided by the proposed registration technique, allowing generalization across categories. Unlike previous work on FSL, training data and support sets contain only positive (normal) examples without any negative (anomalous) samples.
@@ II-C Few-Shot Anomaly Detection
Recent developments in FSAD are shifting towards category-agnostic techniques that enhance feature robustness and generalization without needing parameter fine-tuning. For example, RegAD [27] uses a category-agnostic strategy to perform anomaly detection and localization in new categories. Building on this, RFR [73] uses adversarial loss, common in domain adaptation, to align feature distributions between source and target domains, improving the model’s generalization capabilities. PACKD [74] further enhances generalization by integrating category-specific information from a teacher network into a student network, guiding the learning process with prototypes generated from few-shot normal samples. PromptAD [75] introduces a dual-branch framework that uses prior knowledge about abnormal behaviors through text prompts, leveraging multimodal data to address visual information gaps and enhance AD performance. This paper sets itself apart by introducing a category-agnostic registration mechanism to train the model’s backbone, transforming it into a flexible, plug-and-play tool that bolsters feature robustness and generalization for FSAD. This novel approach signifies a substantial advancement toward more adaptable and broadly applicable AD methods in industrial settings.
@@ IV Category-Agnostic Registration Training
Figure 2: (Left) An overview of the architecture of the proposed category-agnostic registration (CAReg) network. Given a train image and a set of support images, features are extracted by three convolutional residual blocks ($C_{1}$, $C_{2}$, and $C_{3}$), each followed by a spatial transformation module ($S_{1}$, $S_{2}$, and $S_{3}$). A feature registration module is leveraged and supervised by a registration loss (Figure 3). (Right) A spatial transformation module, containing a localization network and a differentiable sampler, is used to learn the mappings, enabling the model to transform features to facilitate feature registration.
@@ IV-A Architecture of the Category-Agnostic Encoder
As shown in Figure 2, a residual convolutional network [79] is used as the feature extractor. Following [22, 12], only the first three convolutional residual blocks of ResNet, i.e., $C_{1}$, $C_{2}$, and $C_{3}$, are kept, while the last convolution block is discarded to ensure the output features retain sufficient spatial information. To enable the alignment of images with different orientation, a spatial transformation module employing a spatial transformer network (STN) [21] is inserted into each convolutional residual block. Specifically, as shown in Figure 2, given a pair of input features $f_{i,a}^{s}$ and $f_{i,B}^{s}=\{f_{i,b}^{s}\}_{b=1}^{k}$ extracted by the convolutional residual block $C_{i}$ for the training image $I_{a}$ and the set of support images $B$, a localization network is leveraged to generate affine transformation matrices $\Theta_{i,a}$ and $\{\Theta_{i,b}\}_{b=1}^{k}$ for feature registration. To perform a warping of the input features, the point-wise transformation is:
@@
Figure 4: Anomaly detection methods and their corresponding memory cost, distribution modeling complexity, and inference complexity. Methods are shown by combining registration trained features extracted by CAReg and three statistical-based normal distribution estimators: (a) PaDim [22], (b) OrthoAD [25], and (c) PatchCore [26]. Symbols used in the complexity equations include: $D$ is the sum of the channel dimensions for the three STN outputs, K is the size of the given support set, $D^{\prime}$ is a constant and $D^{\prime}\ll D$, and $\gamma$ denotes the proportion of the original memory bank that has been sampled. Other symbols are explained in the text.
@@ IV-B Feature Registration Module
Given paired extracted features $f_{3,a}^{t}$ from the training image and $f_{3,B}^{t}=\{f_{3,b}^{t}\}_{b=1}^{k}$ from the support image set as the final transformation outputs, we leverage a Siamese network [80] for feature registration, as shown in Figure 3. The architecture employs a parameter-sharing encoder, $E$, applied on multiple inputs, followed by a prediction head $P$. Training employs the method used by SimSiam [20] to avoid the collapsing problem when optimized without negative pairs. Denoting $p_{a}\triangleq P(E(f_{3,a}^{t}))$ and $z_{B}\triangleq\textit{mean}(E(\{f_{3,b}^{t}\}_{b=1}^{k}))$ where $\textit{mean}(\cdot)$ is an averaging operation on samples, a negative cosine similarity loss is applied:
@@
Different from [27] where only one single image is used as the reference, to improve the robustness, $z_{B}$ and $p_{B}$ are accumulated features (AF) produced by accumulating the features from multiple references. To prevent the collapsing solutions, once the prediction head $P$ is applied on one branch, as shown in Figure 3, a stop-gradient operation is applied to the other branch. Finally, a symmetrized feature registration loss is defined as:
@@ V-A Normal Distribution Estimation
Below we briefly present three popular estimators that may be adopted. These estimators are illustrated in Figure 4.
@@
Complexity Analysis: The complexity analysis of the above three distribution estimator is summarized in Figure 4. It can be seen that compared with PaDim, OrthoAD significantly reduces the memory cost due to its low-rank approximation for the normal distribution $\mathcal{D}_{norm}$. The complexity of the distribution modeling of PatchCore is the minimum modeling complexity among the three estimators.
@@ V-B Augmentation Selection Module
Anomaly detection approaches employ data augmentation as an important tool to expand the dataset, especially in the FSAD scenario [15, 16]. However, the impact of augmentation, specifically, where and what data augmentation to use, has not been fully explored. Instead of simply applying the data augmentations on both the support and test images, this paper shows that augmentations, by expanding the support set, play a very important role in the normal distribution estimation process. We only augment the support set, which also reduces the computational cost. Possible combinations of the augmentations produce a larger augmented support set. We conduct the normal distribution estimation on such an augmented support set.
@@ VI-C Evaluation Protocols and Model Configurations
Evaluation Protocols. The area under the Receiver Operating Characteristic curve metric (AUC) is used to quantify the model performance. For anomaly detection and anomaly localization, AUC at the image-level and the pixel-level are considered, respectively. This metric is widely used for AD performance evaluation.
@@ VI-F2 Spatial Transformation Modules
To investigate the impact of the degree of freedom in the transformation, we conducted experiments with different STN versions, and the results are presented in Table XI in the appendix. Notably, for MVTec, STNs that involve rotation+scale transformations exhibit the best performance. This observation aligns with the fact that most objects in this dataset are centrally located, resulting in similar spatial positions. Consequently, the inclusion of translation in the STNs does not significantly contribute to performance improvement, and the overall effect of the STN modules is limited. In contrast, for MPDD, STNs lead to more substantial performance gains, particularly since large-scale transformations are often required. The images in this dataset exhibit various spatial orientations and positions for objects, making it beneficial to align the features with affine transformations. As a result, the STNs that allow affine transformations achieve the best performance in this scenario.
@@ VI-F4 Data Augmentations and Selection Methods
Figure 6: Visualization of anomaly localization results on MVTec (top three rows) and MPDD (bottom two rows).
@@
Figure 7: The t-SNE visualization of features learned from MVTec. The same t-SNE optimization iterations are used for both methods. CAReg leads to more compacted feature distributions for each category, and larger separation among categories.
Figure 5 displays the selection results for each category obtained Figure 5 shows the selection results for each category using the Wasserstein distance-based method. Table VII indicates that our method, despite using fewer augmentations, outperforms using all augmentations, highlighting the importance of selecting suitable augmentations. Data augmentations have varying effects on objects and textures. For the three objects, performance significantly improves with appropriate augmentations, whereas the selection has little impact on the three textures. This smaller impact on textures may be due to the difficulty in altering their characteristic distribution with simple augmentations. Remarkably, using only four fixed augmentations (graying, flipping, small-angle rotation, small-scale translation) achieves comparable results, especially for tiles, where it performs best.
@@
In addition, we also addressed concerns about possible increased false positive rates from non-independent augmented examples. As shown in Table VIII, thorough evaluations show that our augmentation strategy interestingly reduces the false positive rate (FPR) from 5.44% to 2.65% on MVTec and from 29.3% to 19.8% on MPDD on average with $K=2$ (detailed in Table XII in the appendix). This reduction is largely due to our augmentation selection module, designed to enrich the dataset while avoiding biases or artifacts that could degrade image quality. By carefully selecting beneficial augmentation methods, we maintain dataset integrity and enhance overall model performance.

### v2

@@ Abstract
Most existing anomaly detection (AD) methods require a dedicated model for each category. Such a paradigm, despite its promising results, is computationally expensive and inefficient, thereby failing to meet the requirements for realworld applications. Inspired by how humans detect anomalies, by comparing a query image to known normal ones, this article proposes a novel few-shot AD (FSAD) framework. Using a training set of normal images from various categories, registration, aiming to align normal images of the same categories, is leveraged as the proxy task for self-supervised category-agnostic representation learning. At test time, an image and its corresponding support set, consisting of a few normal images from the same category, are supplied, and anomalies are identified by comparing the registered features of the test image to its corresponding support image features. Such a setup enables the model to generalize to novel test categories. It is, to our best knowledge, the first FSAD method that requires no model fine-tuning for novel categories: enabling a single model to be applied to all categories. Extensive experiments demonstrate the effectiveness of the proposed method. Particularly, it improves the current state-of-the-art (SOTA) for FSAD by 11.3% and 8.3% on the MVTec and MPDD benchmarks, respectively. The source code is available at https://github.com/Haoyan-Guan/CAReg.
@@
Anomaly detection (AD), few-shot learning (FSL), registration, self-supervised learning.
@@ I Introduction
Figure 1: (a) One-model-per-category paradigm for the vanilla AD and FSAD. (b) One-model-all-category paradigm for the proposed category-agnostic FSAD.
@@
Inspired by how human beings detect anomalies, this paper explores a one-model-all-category paradigm for FSAD, i.e., training a category-agnostic AD model. To find the anomalies in an image, humans simply compare it to a normal exemplar and identify any difference. This approach requires no knowledge of the image category because comparison is naturally category agnostic. We thus model AD as a comparison task to make it category agnostic. To deal with the fact that objects may appear in different positions, orientations, or poses, the images are first transformed into one coordinate system through a form of registration [17, 18, 19] to facilitate the comparison.
An overview of the proposed Category-Agnostic Registration (CAReg) framework for FSAD is shown in Fig. 1 (b). A Siamese network [20] containing three spatial transformer network blocks [21] is employed for registration. For better robustness, we perform registration at the feature level, instead of pixel-by-pixel image registration [19]. The feature-level registration loss, a relaxed version of the image-level pixel-wise registration loss, is used to minimize the cosine distance of the features from images of the same category. The trained model has the ability to perform registration across different categories and is able to be used at test-time directly on novel categories. Anomalies are detected by comparing the registered features of the novel test category image and its corresponding support set (a few normal images from the target category). The comparison process first estimates the normal distribution based on the support set, which can be implemented with statistical-based normal distribution estimators [22], and then fits the test samples to the normal distribution. Samples that fall out of the normal distribution are considered anomalies.
@@
  • Feature registration is leveraged as the proxy task to train AD models, which demonstrate increased generalizability across different categories and datasets.
  • CAReg outperforms the state-of-the-art FSAD methods on both AD and anomaly localization tasks with standard benchmark datasets.
@@
The rest of the paper is organized as follows. Sec. II reviews existing works related to AD and few-shot learning. Sec. III formulates the problem of FSAD. Sec. IV describes the category-agnostic registration training used for CAReg. Sec. V describes an AD system combining CAReg with multiple normal distribution estimators. Sec. VI presents the experimental results on several real-world datasets, showing the effectiveness of our method in both FSAD and anomaly localization. Finally, conclusions are drawn in Sec. VII.
@@ II-A Anomaly Detection
AD [28] has a wide range of applications, encompassing still image AD [9, 10, 7, 8] and anomalous human behavior detection [29, 30, 31, 32, 33, 34]. The focus of this paper is on industrial defect detection [23, 24], covering both image-level AD and pixel-level anomaly localization. The main challenge of unsupervised AD, compared to supervised approaches, lies in the absence of training annotations, be it at the image-level and pixel-level. Models are expected to be trained solely with normal instances, excluding anomalies. Unsupervised AD research can be broadly categorized into two tracks: one-class classification based AD approaches and self-supervised learning-based AD approaches.
@@
Recent advances in one-class classification for AD increasingly rely on feature embeddings from pre-trained networks, leveraging these as robust indicators of normalcy. Various statistical estimators of normal distributions [22, 25, 26] help pinpoint significant deviations in feature behavior, flagging them as anomalies. Measures like the Mahalanobis distance [22] are frequently used to calculate anomaly scores. Techniques such as those in [22, 26, 42] have demonstrated notable success by utilizing models pretrained on the ImageNet dataset [43]. However, solely relying on ImageNet pre-training has its limitations. This paper introduces feature registration as a novel proxy task that enhances feature aggregation and generalization capabilities, thus refining AD performance.
@@ II-B Few-Shot Learning
FSL addresses the task of adapting a model to novel categories with limited available images. FSL methods can be broadly categorized into metric learning [64, 65, 66], generation [67, 68, 69], and optimization [70, 71, 72] approaches. Metric learning methods aim to structure the feature space to find the closest distance between test images and known classes. Generation methods involve designing generative models to produce novel class images, thereby bridging the few-shot gap. Optimization methods identify commonalities among different classes and use optimization models for novel classes based on these commonalities. In this paper, CAReg is based on metric learning to address the challenge of FSAD. The challenge is to detect image anomalies for novel categories that are unseen in the training dataset, using only a few normal samples. The key point is to enable the model to be category-agnostic, and this ability is provided by the proposed registration technique, allowing generalization across categories. Unlike previous work on FSL, training data and support sets contain only positive (normal) examples without any negative (anomalous) samples.
@@ II-C Few-Shot Anomaly Detection
Recent developments in FSAD are shifting towards category-agnostic techniques that enhance feature robustness and generalization without needing parameter fine-tuning. For example, RegAD [27] uses a category-agnostic strategy to perform AD and localization in new categories. Building on this, RFR [73] uses adversarial loss, common in domain adaptation, to align feature distributions between source and target domains, improving the model’s generalization capabilities. PACKD [74] further enhances generalization by integrating category-specific information from a teacher network into a student network, guiding the learning process with prototypes generated from few-shot normal samples. PromptAD [75] introduces a dual-branch framework that uses prior knowledge about abnormal behaviors through text prompts, leveraging multimodal data to address visual information gaps and enhance AD performance. This paper sets itself apart by introducing a category-agnostic registration mechanism to train the model’s backbone, transforming it into a flexible, plug-and-play tool that bolsters feature robustness and generalization for FSAD. This novel approach signifies a substantial advancement toward more adaptable and broadly applicable AD methods in industrial settings.
@@ IV Category-Agnostic Registration Training
Figure 2: (Left) An overview of the architecture of the proposed category-agnostic registration (CAReg) network. Given a train image and a set of support images, features are extracted by three convolutional residual blocks ($C_{1}$, $C_{2}$, and $C_{3}$), each followed by a spatial transformation module ($S_{1}$, $S_{2}$, and $S_{3}$). A feature registration module is leveraged and supervised by a registration loss (Fig. 3). (Right) A spatial transformation module, containing a localization network and a differentiable sampler, is used to learn the mappings, enabling the model to transform features to facilitate feature registration.
@@ IV-A Architecture of the Category-Agnostic Encoder
As shown in Fig. 2, a residual convolutional network [79] is used as the feature extractor. Following [22, 12], only the first three convolutional residual blocks of ResNet, i.e., $C_{1}$, $C_{2}$, and $C_{3}$, are kept, while the last convolution block is discarded to ensure the output features retain sufficient spatial information. To enable the alignment of images with different orientation, a spatial transformation module employing a spatial transformer network (STN) [21] is inserted into each convolutional residual block. Specifically, as shown in Fig. 2, given a pair of input features $f_{i,a}^{s}$ and $f_{i,B}^{s}=\{f_{i,b}^{s}\}_{b=1}^{k}$ extracted by the convolutional residual block $C_{i}$ for the training image $I_{a}$ and the set of support images $B$, a localization network is leveraged to generate affine transformation matrices $\Theta_{i,a}$ and $\{\Theta_{i,b}\}_{b=1}^{k}$ for feature registration. To perform a warping of the input features, the point-wise transformation is:
@@
Figure 4: AD methods and their corresponding memory cost, distribution modeling complexity, and inference complexity. Methods are shown by combining registration trained features extracted by CAReg and three statistical-based normal distribution estimators: (a) PaDim [22], (b) OrthoAD [25], and (c) PatchCore [26]. Symbols used in the complexity equations include: $D$ is the sum of the channel dimensions for the three STN outputs, K is the size of the given support set, $D^{\prime}$ is a constant and $D^{\prime}\ll D$, and $\gamma$ denotes the proportion of the original memory bank that has been sampled. Other symbols are explained in the text.
@@ IV-B Feature Registration Module
Given paired extracted features $f_{3,a}^{t}$ from the training image and $f_{3,B}^{t}=\{f_{3,b}^{t}\}_{b=1}^{k}$ from the support image set as the final transformation outputs, we leverage a Siamese network [80] for feature registration, as shown in Fig. 3. The architecture employs a parameter-sharing encoder, $E$, applied on multiple inputs, followed by a prediction head $P$. Training employs the method used by SimSiam [20] to avoid the collapsing problem when optimized without negative pairs. Denoting $p_{a}\triangleq P(E(f_{3,a}^{t}))$ and $z_{B}\triangleq\textit{mean}(E(\{f_{3,b}^{t}\}_{b=1}^{k}))$ where $\textit{mean}(\cdot)$ is an averaging operation on samples, a negative cosine similarity loss is applied:
@@
Different from [27] where only one single image is used as the reference, to improve the robustness, $z_{B}$ and $p_{B}$ are accumulated features (AF) produced by accumulating the features from multiple references. To prevent the collapsing solutions, once the prediction head $P$ is applied on one branch, as shown in Fig. 3, a stop-gradient operation is applied to the other branch. Finally, a symmetrized feature registration loss is defined as:
@@ V-A Normal Distribution Estimation
Below we briefly present three popular estimators that may be adopted. These estimators are illustrated in Fig. 4.
@@
Complexity Analysis: The complexity analysis of the above three distribution estimator is summarized in Fig. 4. It can be seen that compared with PaDim, OrthoAD significantly reduces the memory cost due to its low-rank approximation for the normal distribution $\mathcal{D}_{norm}$. The complexity of the distribution modeling of PatchCore is the minimum modeling complexity among the three estimators.
@@ V-B Augmentation Selection Module
AD approaches employ data augmentation as an important tool to expand the dataset, especially in the FSAD scenario [15, 16]. However, the impact of augmentation, specifically, where and what data augmentation to use, has not been fully explored. Instead of simply applying the data augmentations on both the support and test images, this paper shows that augmentations, by expanding the support set, play a very important role in the normal distribution estimation process. We only augment the support set, which also reduces the computational cost. Possible combinations of the augmentations produce a larger augmented support set. We conduct the normal distribution estimation on such an augmented support set.
@@ VI-C Evaluation Protocols and Model Configurations
Evaluation Protocols. The area under the Receiver Operating Characteristic curve metric (AUC) is used to quantify the model performance. For AD and anomaly localization, AUC at the image-level and the pixel-level are considered, respectively. This metric is widely used for AD performance evaluation.
@@ VI-F2 Spatial Transformation Modules
To investigate the impact of the degree of freedom in the transformation, we conducted experiments with different STN versions, and the results are presented in the appendix. Notably, for MVTec, STNs that involve rotation+scale transformations exhibit the best performance. This observation aligns with the fact that most objects in this dataset are centrally located, resulting in similar spatial positions. Consequently, the inclusion of translation in the STNs does not significantly contribute to performance improvement, and the overall effect of the STN modules is limited. In contrast, for MPDD, STNs lead to more substantial performance gains, particularly since large-scale transformations are often required. The images in this dataset exhibit various spatial orientations and positions for objects, making it beneficial to align the features with affine transformations. As a result, the STNs that allow affine transformations achieve the best performance in this scenario.
@@ VI-F4 Data Augmentations and Selection Methods
Figure 6: Visualization of anomaly localization results on MVTec (top three rows) and MPDD (bottom two rows). (a) Input. PaDim (b) heat map and (c) result. CAReg+PaDim (d) heat map and (e) result. Patchcore (f) heat map and (g) result. CAReg+PatchCore (h) heat map and (i) result. (j) GT.
@@
Figure 7: The t-SNE visualization of features learned from MVTec. The same t-SNE optimization iterations are used for both methods. CAReg leads to more compacted feature distributions for each category, and larger separation among categories. (a) Without feature registration. CAReg.
Fig. 5 displays the selection results for each category obtained Fig. 5 shows the selection results for each category using the Wasserstein distance-based method. Table VII indicates that our method, despite using fewer augmentations, outperforms using all augmentations, highlighting the importance of selecting suitable augmentations. Data augmentations have varying effects on objects and textures. For the three objects, performance significantly improves with appropriate augmentations, whereas the selection has little impact on the three textures. This smaller impact on textures may be due to the difficulty in altering their characteristic distribution with simple augmentations. Remarkably, using only four fixed augmentations (graying, flipping, small-angle rotation, small-scale translation) achieves comparable results, especially for tiles, where it performs best.
@@
In addition, we also addressed concerns about possible increased false positive rates from non-independent augmented examples. As shown in Table VIII, thorough evaluations show that our augmentation strategy interestingly reduces the false positive rate (FPR) from 5.44% to 2.65% on MVTec and from 29.3% to 19.8% on MPDD on average with $K=2$ (detailed in the appendix). This reduction is largely due to our augmentation selection module, designed to enrich the dataset while avoiding biases or artifacts that could degrade image quality. By carefully selecting beneficial augmentation methods, we maintain dataset integrity and enhance overall model performance.